Simplify and express each of the following in exponential form:
 (4^5 × a^8b^3 )/ (4^5 × a^5b^2)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

= 4^5–5× (a^8–5× b^3 - 2) ... [∵a^m ÷a^n = a^m-n ]
= 4^0× (a^3b)
= 1 × a^3b
= a^3b

Thanks